# CultureGuard: Towards Culturally-Aware Dataset and Guard Model for Multilingual Safety Applications

### v1

@@ 5 Conclusion
Leveraging the CultureGuard approach, we curated the Nemotron-Content-Safety-Dataset-Multilingual-v1 (386,661 samples across 9 languages) and used it to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1. This model consistently outperforms major publicly available LLMs on diverse multilingual safety benchmarks, demonstrating a 30.2% improvement over English Llama-Nemotron-Safety-Guard-V2 and 1.3% over multilingual PolyGuard-Qwen, despite the latter’s substantially larger dataset and non-commercial usability. Notably, our model is currently the sole commercially-friendly multilingual safety guard model available.
This work marks a significant step towards bridging the safety gap in multilingual LLMs, enabling the development of truly culturally-aware and robust safety guard models for global deployment. The dataset and model will be publicly released.
@@
Our work is grounded in ethical considerations, beginning with the use of source datasets and models exclusively under commercial-friendly licenses. The CultureGuard dataset and the resulting model will be released to the research community under a commercial-permissive license to foster further innovation in LLM safety. However, by its very nature, the dataset contains critically unsafe and offensive content intended for training robust safety systems. We urge users to exercise extreme caution and handle the data responsibly.
@@ Appendix A Appendix
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 63.92 86.26 56.53 65.84 69.15 64.44 59.03 50.00 58.38 65.63 Llama-Guard-3-8B 71.78 77.35 70.02 72.42 72.96 72.86 70.91 70.59 68.04 70.91 Llama-Guard-4-12B 67.47 71.64 65.52 66.86 66.90 67.33 64.64 70.92 65.51 67.88 PolyGuard-Qwen 84.94 86.26 84.51 85.95 85.53 85.86 82.06 85.11 83.48 85.69 CultureGuard (vanilla) 84.84 86.58 84.09 85.44 85.35 85.91 83.46 85.17 83.27 84.34 CultureGuard 85.58 86.77 84.20 85.84 85.63 86.04 84.55 86.30 84.82 86.04 CultureGuard + JB SDG 85.43 86.67 84.93 85.98 85.29 86.30 84.00 86.29 84.01 85.42 CultureGuard + JB SDG + QF 85.15 86.50 84.77 85.33 84.92 85.46 84.12 85.44 84.07 85.77
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 78.96 87.34 69.29 81.52 80.74 82.60 79.66 71.11 78.16 80.17 Llama-Guard-3-8B 60.54 64.34 58.46 60.30 61.59 60.10 61.99 60.47 58.92 58.70 Llama-Guard-4-12B 57.79 64.18 57.19 60.27 58.82 58.33 55.67 56.74 53.97 54.96 PolyGuard-Qwen 80.41 80.89 81.11 80.69 79.00 80.00 80.05 80.79 80.47 80.73 CultureGuard (vanilla) 85.46 87.67 83.94 85.23 86.27 85.92 85.89 84.62 84.58 85.05 CultureGuard 85.56 86.93 84.75 85.26 86.45 84.89 85.51 86.18 84.07 85.96 CultureGuard + JB SDG 84.57 86.42 82.78 84.36 86.17 83.90 85.37 84.44 83.45 84.24 CultureGuard + JB SDG + QF 85.48 86.36 85.30 83.96 86.56 86.12 84.72 86.20 84.95 85.15
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.58 88.51 61.72 69.05 79.02 73.35 55.65 42.86 62.35 66.67 Llama-Guard-3-8B 81.15 83.71 82.23 82.57 83.29 85.17 79.33 80.93 73.32 79.80 Llama-Guard-4-12B 74.79 77.39 72.53 74.45 75.00 75.26 69.43 81.52 72.24 75.32 PolyGuard-Qwen 95.13 88.77 96.05 97.21 96.49 98.31 92.86 95.82 94.07 96.60 CultureGuard (vanilla) 92.97 88.41 93.75 93.50 93.53 95.69 93.25 94.09 91.36 93.13 CultureGuard 94.84 89.17 95.63 95.62 95.04 96.83 95.53 96.00 94.78 94.99 CultureGuard + JB SDG 94.99 89.08 96.75 95.20 95.88 97.24 94.82 96.25 94.95 94.74 CultureGuard + JB SDG + QF 95.11 89.65 96.34 94.49 96.37 97.06 94.57 97.18 94.81 95.52
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 82.87 81.16 70.97 91.23 92.06 91.53 81.19 68.18 87.76 81.72 Llama-Guard-3-8B 60.18 58.23 59.34 62.50 65.42 57.45 58.43 62.22 57.83 60.24 Llama-Guard-4-12B 59.11 55.84 60.67 66.67 61.54 63.92 54.12 55.42 59.26 54.55 PolyGuard-Qwen 85.72 76.68 87.88 89.05 87.18 84.72 88.52 83.97 89.08 84.38 CultureGuard (vanilla) 92.68 83.50 92.31 95.24 95.59 96.00 90.27 92.98 94.64 93.58 CultureGuard 93.32 82.13 92.98 96.00 94.74 95.24 94.02 94.74 96.43 93.58 CultureGuard + JB SDG 93.63 80.18 93.33 94.49 98.53 96.12 93.22 96.55 95.65 94.55 CultureGuard + JB SDG + QF 93.52 81.06 94.31 95.31 96.40 96.06 93.33 97.44 95.58 92.17
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 57.62 82.69 52.17 56.58 61.01 60.07 52.57 41.76 52.97 58.77 Llama-Guard-3-8B 67.34 76.18 63.33 68.05 69.93 69.74 66.72 63.25 61.76 67.10 Llama-Guard-4-12B 62.02 73.82 58.02 60.34 65.21 62.60 59.64 59.21 57.77 61.59 PolyGuard-Qwen 84.68 87.54 84.42 84.33 85.37 84.30 83.02 84.81 83.47 84.85 CultureGuard (vanilla) 74.85 81.08 72.88 73.71 77.12 75.89 72.34 72.89 72.68 75.02 CultureGuard 78.88 84.82 77.02 78.62 80.49 79.51 75.93 77.03 76.96 79.56 CultureGuard + JB SDG 79.35 83.78 78.31 78.64 80.54 79.47 78.27 78.21 77.70 79.27 CultureGuard + JB SDG + QF 79.25 83.91 77.74 78.91 81.15 79.28 77.73 77.28 77.79 79.44
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 68.50 75.79 63.30 71.67 72.29 70.72 64.68 62.20 67.28 68.52 Llama-Guard-3-8B 64.68 69.62 62.71 65.59 66.53 66.53 62.87 65.29 60.49 62.53 Llama-Guard-4-12B 53.08 66.13 46.26 55.75 55.08 56.11 44.60 51.28 48.93 53.55 PolyGuard-Qwen 74.19 77.68 77.51 73.54 71.21 73.68 74.87 73.27 75.09 70.83 CultureGuard (vanilla) 71.38 76.41 70.23 70.86 73.40 73.46 68.40 70.61 69.17 69.85 CultureGuard 70.82 73.14 69.83 71.72 73.65 72.50 66.67 70.15 68.43 71.27 CultureGuard + JB SDG 73.44 76.29 74.62 72.73 74.04 73.40 71.53 72.43 71.28 74.66 CultureGuard + JB SDG + QF 72.89 74.28 73.65 72.55 73.40 73.28 72.32 71.95 71.59 73.02
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.45 97.27 24.54 74.74 83.05 79.49 59.43 58.43 43.64 77.46 Llama-Guard-3-8B 47.01 49.45 44.19 50.42 49.19 48.99 51.97 46.27 36.78 45.80 Llama-Guard-4-12B 39.12 38.75 43.95 42.25 33.61 39.17 38.39 51.47 29.35 35.18 PolyGuard-Qwen 86.98 91.60 82.76 90.40 89.67 90.40 80.46 89.54 76.73 91.28 CultureGuard (vanilla) 90.39 97.00 85.85 91.08 92.08 93.34 87.49 91.68 81.16 93.79 CultureGuard 89.87 97.05 84.17 90.82 91.88 93.75 86.89 92.27 78.53 93.48 CultureGuard + JB SDG 89.97 96.42 84.27 91.01 91.69 93.43 87.17 92.11 79.44 94.23 CultureGuard + JB SDG + QF 91.49 96.17 88.94 92.29 91.59 93.29 90.29 92.53 83.95 94.40
@@
Models Average en ar th zh Nemotron-Safety-Guard-V2 72.08 93.48 60.62 63.34 70.90 Llama-Guard-3-8B 78.04 79.92 76.95 79.54 75.74 Llama-Guard-4-12B 73.56 76.92 67.09 74.50 75.74 PolyGuard-Qwen 93.67 94.90 91.75 93.04 95.00 CultureGuard (vanilla) 92.91 93.43 92.49 93.04 92.67 CultureGuard 93.69 94.58 92.67 93.94 93.58 CultureGuard + JB SDG 93.94 95.20 92.31 93.94 94.30 CultureGuard + JB SDG + QF 95.36 95.58 94.30 95.35 96.21
@@
Models Average en ar de fr hi ja zh Nemotron-Safety-Guard-V2 35.31 74.96 25.16 30.92 33.43 23.33 20.17 39.22 Llama-Guard-3-8B 40.82 58.47 36.97 37.11 38.99 36.40 37.78 40.00 Llama-Guard-4-12B 39.13 54.11 40.64 37.63 38.24 31.34 37.11 34.81 PolyGuard-Qwen 64.50 77.22 62.24 61.52 62.58 60.79 64.08 63.05 CultureGuard (vanilla) 59.88 75.14 58.30 56.08 57.40 57.23 57.40 57.65 CultureGuard 62.90 75.47 60.87 58.51 59.15 62.70 62.48 61.15 CultureGuard + JB SDG 62.77 75.63 60.17 59.61 60.35 59.61 61.52 62.48 CultureGuard + JB SDG + QF 66.97 75.86 66.44 64.24 63.68 65.73 66.63 66.22
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.73 91.39 55.82 68.78 74.73 68.49 59.33 47.40 63.70 70.95 Llama-Guard-3-8B 72.29 80.92 63.03 74.24 77.11 74.42 70.24 69.17 67.82 73.67 Llama-Guard-4-12B 69.63 82.00 60.61 73.12 73.53 70.29 64.21 66.53 63.15 73.22 PolyGuard-Qwen 88.97 93.14 83.26 90.39 92.53 90.06 88.28 86.02 86.91 90.14 CultureGuard (vanilla) 86.78 92.80 80.50 87.16 90.88 88.50 86.12 82.84 84.36 87.88 CultureGuard 88.53 94.68 82.58 90.16 91.29 88.87 88.67 84.93 85.69 89.88 CultureGuard + JB SDG 91.54 96.35 85.05 92.97 93.90 91.22 92.57 87.70 90.97 93.16 CultureGuard + JB SDG + QF 91.77 95.57 85.13 93.37 93.70 91.84 92.27 89.11 90.36 94.59
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 86.55 96.11 80.68 87.82 89.09 89.61 86.70 79.41 81.50 88.02 Llama-Guard-3-8B 77.66 82.30 75.59 79.70 81.19 80.30 75.39 76.02 72.58 75.84 Llama-Guard-4-12B 70.14 79.63 65.72 74.75 74.87 68.65 65.14 65.00 64.59 72.92 PolyGuard-Qwen 89.89 93.50 90.99 88.66 86.71 89.88 88.79 89.84 88.61 92.08 CultureGuard (vanilla) 91.23 92.96 90.02 91.02 92.50 92.44 90.79 90.95 87.85 92.51 CultureGuard 91.99 95.18 90.38 92.15 93.42 93.02 90.35 90.64 88.89 93.87 CultureGuard + JB SDG 94.64 96.77 93.74 95.54 95.93 95.50 93.45 93.89 93.16 93.79 CultureGuard + JB SDG + QF 94.35 96.61 94.27 94.65 95.36 95.32 93.39 93.72 92.60 93.28

### v2

@@ 5 Conclusion
Leveraging the CultureGuard approach, we curated the Nemotron-Content-Safety-Dataset-Multilingual-v1 (386,661 samples across 9 languages) and used it to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1. This model consistently outperforms major publicly available LLMs on diverse multilingual safety benchmarks, demonstrating a 31.06% improvement over English Llama-Nemotron-Safety-Guard-V2 and 1.28% over multilingual PolyGuard-Qwen, despite the latter’s substantially larger dataset and non-commercial usability. Notably, our model is currently the sole commercially-friendly multilingual safety guard model available.
This work marks a significant step towards bridging the safety gap in multilingual LLMs, enabling the development of truly culturally-aware and robust safety guard models for global deployment.
Limitations
Our adoption of a fully synthetic pipeline offers significant advantages in scalability and efficiency, enabling the rapid generation of large-scale data. This approach, however, comes with certain trade-offs. The quality of the generated data is naturally linked to the capabilities of the underlying LLMs, and the fidelity of the final dataset is influenced by the accuracy of machine translation and the efficacy of our automated filtering methods. Additionally, since the cultural adaptation operates on English text with a fixed safety taxonomy, some deep linguistic nuances or novel, culture-specific harm categories may not be fully captured. We acknowledge these aspects as opportunities for refinement in future iterations.
@@
Our work is grounded in ethical considerations, beginning with the use of source datasets and models exclusively under commercial-friendly licenses. The CultureGuard dataset and the resulting model is released to the research community under a commercial-permissive license to foster further innovation in LLM safety. However, by its very nature, the dataset contains critically unsafe and offensive content intended for training robust safety systems. We urge users to exercise extreme caution and handle the data responsibly.
@@ Appendix A Appendix
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 63.92 86.26 56.53 65.84 69.15 64.44 59.03 50.00 58.38 65.63 Llama-Guard-3-8B 71.78 77.35 70.02 72.42 72.96 72.86 70.91 70.59 68.04 70.91 Llama-Guard-4-12B 67.47 71.64 65.52 66.86 66.90 67.33 64.64 70.92 65.51 67.88 Granite Guardian 3.1 8B 78.63 83.71 75.52 81.97 81.64 82.71 73.86 80.80 66.23 81.24 Granite Guardian 3.3 8B (reasoning off) 75.43 85.21 72.61 81.80 81.99 82.68 70.12 77.82 46.32 80.28 Granite Guardian 3.3 8B (reasoning on) 76.87 84.24 73.46 80.70 80.88 81.35 74.64 76.93 60.01 79.60 PolyGuard-Qwen 84.94 86.26 84.51 85.95 85.53 85.86 82.06 85.11 83.48 85.69 CultureGuard (vanilla) 84.84 86.58 84.09 85.44 85.35 85.91 83.46 85.17 83.27 84.34 CultureGuard 85.58 86.77 84.20 85.84 85.63 86.04 84.55 86.30 84.82 86.04 CultureGuard + JB SDG 85.43 86.67 84.93 85.98 85.29 86.30 84.00 86.29 84.01 85.42 CultureGuard + JB SDG + QF 85.15 86.50 84.77 85.33 84.92 85.46 84.12 85.44 84.07 85.77
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 78.96 87.34 69.29 81.52 80.74 82.60 79.66 71.11 78.16 80.17 Llama-Guard-3-8B 60.54 64.34 58.46 60.30 61.59 60.10 61.99 60.47 58.92 58.70 Llama-Guard-4-12B 57.79 64.18 57.19 60.27 58.82 58.33 55.67 56.74 53.97 54.96 Granite Guardian 3.1 8B 75.16 78.07 74.32 77.35 78.24 78.05 72.06 77.17 62.17 79.02 Granite Guardian 3.3 8B (reasoning off) 75.86 80.62 74.57 79.45 80.35 81.27 73.11 77.63 55.50 80.21 Granite Guardian 3.3 8B (reasoning on) 72.27 77.45 67.75 75.00 76.33 77.26 70.16 74.27 56.96 75.23 PolyGuard-Qwen 80.41 80.89 81.11 80.69 79.00 80.00 80.05 80.79 80.47 80.73 CultureGuard (vanilla) 85.46 87.67 83.94 85.23 86.27 85.92 85.89 84.62 84.58 85.05 CultureGuard 85.56 86.93 84.75 85.26 86.45 84.89 85.51 86.18 84.07 85.96 CultureGuard + JB SDG 84.57 86.42 82.78 84.36 86.17 83.90 85.37 84.44 83.45 84.24 CultureGuard + JB SDG + QF 85.48 86.36 85.30 83.96 86.56 86.12 84.72 86.20 84.95 85.15
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.58 88.51 61.72 69.05 79.02 73.35 55.65 42.86 62.35 66.67 Llama-Guard-3-8B 81.15 83.71 82.23 82.57 83.29 85.17 79.33 80.93 73.32 79.80 Llama-Guard-4-12B 74.79 77.39 72.53 74.45 75.00 75.26 69.43 81.52 72.24 75.32 Granite Guardian 3.1 8B 85.43 86.43 83.79 90.25 90.59 92.65 79.07 90.02 65.35 90.70 Granite Guardian 3.3 8B (reasoning off) 81.14 87.03 80.72 89.22 88.94 91.44 70.80 86.96 47.06 88.11 Granite Guardian 3.3 8B (reasoning on) 80.13 86.08 78.24 84.44 87.05 88.68 73.05 80.10 57.66 85.85 PolyGuard-Qwen 95.13 88.77 96.05 97.21 96.49 98.31 92.86 95.82 94.07 96.60 CultureGuard (vanilla) 92.97 88.41 93.75 93.50 93.53 95.69 93.25 94.09 91.36 93.13 CultureGuard 94.84 89.17 95.63 95.62 95.04 96.83 95.53 96.00 94.78 94.99 CultureGuard + JB SDG 94.99 89.08 96.75 95.20 95.88 97.24 94.82 96.25 94.95 94.74 CultureGuard + JB SDG + QF 95.11 89.65 96.34 94.49 96.37 97.06 94.57 97.18 94.81 95.52
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 82.87 81.16 70.97 91.23 92.06 91.53 81.19 68.18 87.76 81.72 Llama-Guard-3-8B 60.18 58.23 59.34 62.50 65.42 57.45 58.43 62.22 57.83 60.24 Llama-Guard-4-12B 59.11 55.84 60.67 66.67 61.54 63.92 54.12 55.42 59.26 54.55 Granite Guardian 3.1 8B 75.79 68.90 73.79 85.96 85.04 86.67 63.16 82.57 53.16 82.83 Granite Guardian 3.3 8B (reasoning off) 81.87 75.11 79.63 91.20 88.89 92.80 68.82 87.04 63.41 89.91 Granite Guardian 3.3 8B (reasoning on) 76.93 70.37 73.27 82.14 83.72 89.26 68.09 76.64 65.12 83.81 PolyGuard-Qwen 85.72 76.68 87.88 89.05 87.18 84.72 88.52 83.97 89.08 84.38 CultureGuard (vanilla) 92.68 83.50 92.31 95.24 95.59 96.00 90.27 92.98 94.64 93.58 CultureGuard 93.32 82.13 92.98 96.00 94.74 95.24 94.02 94.74 96.43 93.58 CultureGuard + JB SDG 93.63 80.18 93.33 94.49 98.53 96.12 93.22 96.55 95.65 94.55 CultureGuard + JB SDG + QF 93.52 81.06 94.31 95.31 96.40 96.06 93.33 97.44 95.58 92.17
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 57.62 82.69 52.17 56.58 61.01 60.07 52.57 41.76 52.97 58.77 Llama-Guard-3-8B 67.34 76.18 63.33 68.05 69.93 69.74 66.72 63.25 61.76 67.10 Llama-Guard-4-12B 62.02 73.82 58.02 60.34 65.21 62.60 59.64 59.21 57.77 61.59 Granite Guardian 3.1 8B 76.87 84.18 76.57 80.00 81.05 80.35 71.46 77.42 60.61 80.17 Granite Guardian 3.3 8B (reasoning off) 72.30 84.35 70.87 76.70 78.63 76.99 69.34 71.51 44.25 78.10 Granite Guardian 3.3 8B (reasoning on) 68.47 80.06 66.12 70.26 72.97 72.09 65.95 65.95 51.46 71.38 PolyGuard-Qwen 84.68 87.54 84.42 84.33 85.37 84.30 83.02 84.81 83.47 84.85 CultureGuard (vanilla) 74.85 81.08 72.88 73.71 77.12 75.89 72.34 72.89 72.68 75.02 CultureGuard 78.88 84.82 77.02 78.62 80.49 79.51 75.93 77.03 76.96 79.56 CultureGuard + JB SDG 79.35 83.78 78.31 78.64 80.54 79.47 78.27 78.21 77.70 79.27 CultureGuard + JB SDG + QF 79.25 83.91 77.74 78.91 81.15 79.28 77.73 77.28 77.79 79.44
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 68.50 75.79 63.30 71.67 72.29 70.72 64.68 62.20 67.28 68.52 Llama-Guard-3-8B 64.68 69.62 62.71 65.59 66.53 66.53 62.87 65.29 60.49 62.53 Llama-Guard-4-12B 53.08 66.13 46.26 55.75 55.08 56.11 44.60 51.28 48.93 53.55 Granite Guardian 3.1 8B 66.37 74.29 66.00 71.29 71.14 69.87 56.73 70.39 46.08 71.53 Granite Guardian 3.3 8B (reasoning off) 67.92 75.05 64.95 72.40 72.73 71.83 65.36 68.19 47.94 72.80 Granite Guardian 3.3 8B (reasoning on) 61.22 69.70 59.82 61.11 67.76 65.02 57.14 59.23 47.03 64.15 PolyGuard-Qwen 74.19 77.68 77.51 73.54 71.21 73.68 74.87 73.27 75.09 70.83 CultureGuard (vanilla) 71.38 76.41 70.23 70.86 73.40 73.46 68.40 70.61 69.17 69.85 CultureGuard 70.82 73.14 69.83 71.72 73.65 72.50 66.67 70.15 68.43 71.27 CultureGuard + JB SDG 73.44 76.29 74.62 72.73 74.04 73.40 71.53 72.43 71.28 74.66 CultureGuard + JB SDG + QF 72.89 74.28 73.65 72.55 73.40 73.28 72.32 71.95 71.59 73.02
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.45 97.27 24.54 74.74 83.05 79.49 59.43 58.43 43.64 77.46 Llama-Guard-3-8B 47.01 49.45 44.19 50.42 49.19 48.99 51.97 46.27 36.78 45.80 Llama-Guard-4-12B 39.12 38.75 43.95 42.25 33.61 39.17 38.39 51.47 29.35 35.18 Granite Guardian 3.1 8B 86.26 91.88 80.08 88.43 89.79 91.50 85.92 86.77 71.13 90.83 Granite Guardian 3.3 8B (reasoning off) 78.23 94.35 66.00 88.12 88.34 91.33 59.59 79.07 47.69 89.62 Granite Guardian 3.3 8B (reasoning on) 78.72 91.14 68.73 86.71 88.39 90.03 66.62 81.75 46.75 88.33 PolyGuard-Qwen 86.98 91.60 82.76 90.40 89.67 90.40 80.46 89.54 76.73 91.28 CultureGuard (vanilla) 90.39 97.00 85.85 91.08 92.08 93.34 87.49 91.68 81.16 93.79 CultureGuard 89.87 97.05 84.17 90.82 91.88 93.75 86.89 92.27 78.53 93.48 CultureGuard + JB SDG 89.97 96.42 84.27 91.01 91.69 93.43 87.17 92.11 79.44 94.23 CultureGuard + JB SDG + QF 91.49 96.17 88.94 92.29 91.59 93.29 90.29 92.53 83.95 94.40
@@
Models Average en ar th zh Nemotron-Safety-Guard-V2 72.08 93.48 60.62 63.34 70.90 Llama-Guard-3-8B 78.04 79.92 76.95 79.54 75.74 Llama-Guard-4-12B 73.56 76.92 67.09 74.50 75.74 Granite Guardian 3.1 8B 87.21 94.90 84.62 77.19 92.12 Granite Guardian 3.3 8B (reasoning off) 80.99 94.89 80.46 56.49 92.12 Granite Guardian 3.3 8B (reasoning on) 85.71 94.53 84.40 71.95 91.94 PolyGuard-Qwen 93.67 94.90 91.75 93.04 95.00 CultureGuard (vanilla) 92.91 93.43 92.49 93.04 92.67 CultureGuard 93.69 94.58 92.67 93.94 93.58 CultureGuard + JB SDG 93.94 95.20 92.31 93.94 94.30 CultureGuard + JB SDG + QF 95.36 95.58 94.30 95.35 96.21
@@
Models Average en ar de fr hi ja zh Nemotron-Safety-Guard-V2 35.31 74.96 25.16 30.92 33.43 23.33 20.17 39.22 Llama-Guard-3-8B 40.82 58.47 36.97 37.11 38.99 36.40 37.78 40.00 Llama-Guard-4-12B 39.13 54.11 40.64 37.63 38.24 31.34 37.11 34.81 Granite Guardian 3.1 8B 64.98 73.14 59.33 60.17 63.78 68.13 64.24 66.06 Granite Guardian 3.3 8B (reasoning off) 56.93 71.18 48.16 56.63 59.58 48.58 54.47 59.89 Granite Guardian 3.3 8B (reasoning on) 58.91 70.01 54.05 57.40 59.79 53.75 56.04 61.32 PolyGuard-Qwen 64.50 77.22 62.24 61.52 62.58 60.79 64.08 63.05 CultureGuard (vanilla) 59.88 75.14 58.30 56.08 57.40 57.23 57.40 57.65 CultureGuard 62.90 75.47 60.87 58.51 59.15 62.70 62.48 61.15 CultureGuard + JB SDG 62.77 75.63 60.17 59.61 60.35 59.61 61.52 62.48 CultureGuard + JB SDG + QF 66.97 75.86 66.44 64.24 63.68 65.73 66.63 66.22
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 66.73 91.39 55.82 68.78 74.73 68.49 59.33 47.40 63.70 70.95 Llama-Guard-3-8B 72.29 80.92 63.03 74.24 77.11 74.42 70.24 69.17 67.82 73.67 Llama-Guard-4-12B 69.63 82.00 60.61 73.12 73.53 70.29 64.21 66.53 63.15 73.22 Granite Guardian 3.1 8B 83.59 88.44 79.15 86.64 89.62 86.80 80.32 83.63 70.74 87.02 Granite Guardian 3.3 8B (reasoning off) 82.40 93.79 75.22 88.01 92.11 89.16 81.65 81.72 50.44 89.53 Granite Guardian 3.3 8B (reasoning on) 81.37 92.72 72.37 85.77 88.51 86.31 80.80 78.29 61.02 86.53 PolyGuard-Qwen 88.97 93.14 83.26 90.39 92.53 90.06 88.28 86.02 86.91 90.14 CultureGuard (vanilla) 86.78 92.80 80.50 87.16 90.88 88.50 86.12 82.84 84.36 87.88 CultureGuard 88.53 94.68 82.58 90.16 91.29 88.87 88.67 84.93 85.69 89.88 CultureGuard + JB SDG 91.54 96.35 85.05 92.97 93.90 91.22 92.57 87.70 90.97 93.16 CultureGuard + JB SDG + QF 91.77 95.57 85.13 93.37 93.70 91.84 92.27 89.11 90.36 94.59
@@
Models Average en ar de es fr hi ja th zh Nemotron-Safety-Guard-V2 86.55 96.11 80.68 87.82 89.09 89.61 86.70 79.41 81.50 88.02 Llama-Guard-3-8B 77.66 82.30 75.59 79.70 81.19 80.30 75.39 76.02 72.58 75.84 Llama-Guard-4-12B 70.14 79.63 65.72 74.75 74.87 68.65 65.14 65.00 64.59 72.92 Granite Guardian 3.1 8B 80.93 91.86 78.09 81.86 86.90 84.24 72.78 82.66 62.68 87.36 Granite Guardian 3.3 8B (reasoning off) 83.97 91.40 81.59 86.05 89.50 89.75 80.71 81.28 68.51 86.92 Granite Guardian 3.3 8B (reasoning on) 78.55 84.10 79.20 80.88 82.33 82.10 74.67 76.19 68.65 78.80 PolyGuard-Qwen 89.89 93.50 90.99 88.66 86.71 89.88 88.79 89.84 88.61 92.08 CultureGuard (vanilla) 91.23 92.96 90.02 91.02 92.50 92.44 90.79 90.95 87.85 92.51 CultureGuard 91.99 95.18 90.38 92.15 93.42 93.02 90.35 90.64 88.89 93.87 CultureGuard + JB SDG 94.64 96.77 93.74 95.54 95.93 95.50 93.45 93.89 93.16 93.79 CultureGuard + JB SDG + QF 94.35 96.61 94.27 94.65 95.36 95.32 93.39 93.72 92.60 93.28
@@
Models Average en ar es fr hi Nemotron-Safety-Guard-V2 70.95 89.78 43.61 77.53 72.77 71.04 Llama-Guard-3-8B 80.89 76.93 73.15 84.06 83.61 86.69 Llama-Guard-4-12B 70.01 63.11 65.97 73.36 72.17 75.43 Granite Guardian 3.1 8B 93.79 95.72 88.61 96.42 95.77 92.42 Granite Guardian 3.3 8B (reasoning off) 91.08 96.05 78.79 96.36 95.84 88.35 Granite Guardian 3.3 8B (reasoning on) 92.17 96.32 81.81 95.52 96.30 90.88 PolyGuard-Qwen 95.46 95.17 93.43 97.91 96.37 94.40 CultureGuard (vanilla) 93.81 89.60 92.34 96.36 94.42 96.32 CultureGuard 94.90 91.84 92.09 97.91 95.43 97.25 CultureGuard + JB SDG 95.52 93.41 93.93 97.64 95.77 96.84 CultureGuard + JB SDG + QF 96.79 94.83 96.73 98.24 96.63 97.54
Table 16: Harmful content classification performance (f1-score) on Aya Red-teaming test set (prompt classification).